OpenStudy (anonymous):

solve the following system of equations. 4x-3y=5 2x+2y=6

OpenStudy (anonymous):

x=2 y=1

OpenStudy (diyadiya):

4x-3y=5-- (1) 2x+2y=6--(2) multiply (2) by 2 4x-3y=5--(3) 4x+4y=12--(4) (4)-(3) 4x + 4y =12 4x - 3y = 5 ----------- 0 7y = 7 y=1 4x - 3(1)=5 4x=5+3 x=2 x=2 y=1